# Test-Time Training Provably Improves Transformers as In-context Learners

### v1

@@ 1 Introduction
Modern language models can be viewed as general-purpose, addressing a diverse range of user queries in a zero-shot fashion (Kojima et al., 2022; Wei et al., 2022). However, as the complexity or novelty of the query increases, such as in complex multi-step reasoning scenarios, the pretrained model may falter. This has motivated two popular approaches: in-context learning and test-time computation. In-context learning (ICL) incorporates demonstrations related to the query as part of the prompt facilitating better inference by the model (Brown et al., 2020; Min et al., 2022). Test-time computation methods explicitly increase the inference-time compute to elicit higher quality responses (Snell et al., 2024; Jaech et al., 2024). An important instance of test-time computation is test-time training (TTT) where the weights of a model are explicitly updated to adapt to specific test instances (Sun et al., 2020; Liu et al., 2021). A gradient-based TTT approach can be described as follows: Given a test prompt and a pretrained sequence model, we update the pretrained weights by performing a few gradient iterations on a suitable self-supervised objective (e.g. next-token prediction objective). We can then use the resulting model to perform the inference on the test prompt. This procedure is applicable beyond language models, and it is conceptually similar to meta-learning approaches such as model-agnostic meta-learning (Finn et al., 2017).
Recently, TTT has found significant success in the context of language modeling by boosting the accuracy and reasoning capability of state-of-the-art models (Akyürek et al., 2024; Sun et al., 2024). This success is in part due to the fact that TTT can be naturally integrated within in-context learning: We can fine-tune the transformer model to fit to the labels or chain-of-thought rationales provided as part of the in-context demonstrations, and then re-apply ICL on the adapted model. In fact, recent work (Akyürek et al., 2024) utilizes a variation of this procedure and additional data augmentation to obtain remarkable improvements in the ARC reasoning benchmark. This motivates the central question of our work:
@@
Contributions. As our central contribution, we address this question by providing a comprehensive theoretical study of test-time training for one-layer linear transformers. Focusing on prompts following a linear dataset model, we provide a precise risk characterization of TTT with one gradient step update rule. This characterization is established in terms of three ingredients: (i) context length (number of in-context examples during inference), (ii) target sample size available for TTT in Section 4, and (iii) alignment between the pre-trained model and target task in Section 5. We show that, as sample size increases, TTT can alleviate the distribution shift bottlenecks that arise in standard ICL. This also reveals regimes where TTT with zero or small initialization is preferable to TTT with the pre-trained model (i.e. cold vs. warm start). Interestingly, our experiments with the GPT2 architecture (Radford et al., 2019) show that multilayer transformers exhibit behavior in line with our theory. Additionally, while standard ICL requires $\Omega(d)$ context length under an isotropic task prior, with $d$ denoting feature dimension, we prove that TTT can succeed with $o(d)$ context length by effectively memorizing the target task. Our technical novelty stems from accurately capturing the statistical benefit of TTT during in-context learning. Specifically, while the transformer is trained on a single prompt, we characterize how the sample complexity benefit of TTT is proportional to the number of target examples within the prompt.
As empirical corroboration of our theory, we explore tabular learning and TabPFN (Hollmann et al., 2023, 2025) – a state-of-the-art tabular foundation model pretrained with structural causal model priors. TabPFN is well-aligned with our theoretical setting with similar token encodings but different prior distributions. An important drawback of TabPFN lies in its inference cost, as it uses a full tabular dataset as context during inference. In line with our theory, we demonstrate that TTT can convert TabPFN into a task-specific tabular model that works equally well with significantly less data (up to 5 times fewer). This in turn implies substantial inference gains given that the complexity of softmax-attention is quadratic in the sequence length.
@@ 2 Related Work
In-context learning. In-context learning has received significant interest during the past few years (Brown et al., 2020; Liu et al., 2023; Agarwal et al., 2024). This has also motivated works toward stronger theoretical understanding of ICL (Zhang et al., 2024; Mahankali et al., 2024; Ahn et al., 2023b; Xie et al., 2022; Garg et al., 2022; Li et al., 2023). Closer to us, Mahankali et al. (2024); Ahn et al. (2023b); Zhang et al. (2024) study the optimization landscape of one-layer linear attention model and show that the optimized model implements a single step of projected gradient descent over the in-context demonstrations. More recent works (Lu et al., 2024; Wu et al., 2023) extend these to characterize the pretraining task sample complexity of ICL. While these works focus on pretraining capabilities, we focus on adaptation of the pretrained model to the target task. Additionally, rather than empirical risk minimization, we use gradient descent for adaptation and characterize its risk when the sample size is determined by the target prompt length. In our theoretical model, each token represents a datapoint containing input features and the target label. Remarkably, TabPFN (Hollmann et al., 2023, 2025) shows that, using this simple encoding and pretraining the model with sufficiently rich data priors, transformers can accomplish state-of-the-art classification on tabular datasets via in-context learning.
Test-time training. Test-time training (Sun et al., 2020; Liu et al., 2021; Gandelsman et al., 2022) and related test-time adaptation (Wang et al., 2021; Niu et al., 2022; Yuan et al., 2023) methods aim to overcome distribution shift bottlenecks. This is typically accomplished by adapting the model with the test example using self-supervised or unsupervised objectives. These methods can work with just a single text example or admit a minibatch of examples. For sequence/language modeling tasks, one can utilize TTT on the test sequence via the next-token prediction objective (Sun et al., 2024). Specifically, during in-context learning, the query is unlabeled (e.g. math problem to solve), but we are provided with related examples and associated labels (e.g. through retrieval). Thus, we can fit to these labels as the source of supervision (essentially fine-tuning the model on the dataset of in-context examples). For instance, the training objective in Akyürek et al. (2024) utilizes this approach boosted by additional data augmentation. Finally, the computational efficiency of TTT is an important consideration which motivates our investigation of TTT with a single gradient update.
@@ 3 Problem Setup
Notation. Let $[n]$ denote the set $\{1,\cdots,n\}$ for an integer $n\geq 1$. We denote vectors and matrices using bold lower-case and upper-case letters, respectively, such as $\bm{x}$ for vectors and ${\bm{X}}$ for matrices. The element $x_{i}$ refers to the $i$-th entry of the vector $\bm{x}$. We represent the zero vector of size $n$ by $0_{n}$ and the zero matrix of size $m\times n$ by $\bm{0}_{m\times n}$. The operator $\texttt{tr}({\bm{X}})$ represents the trace of ${\bm{X}}$, ${\bm{X}}^{\dagger}$ is its Moore–Penrose pseudoinverse, and $\|{\bm{X}}\|_{2}$ denotes the spectral norm of ${\bm{X}}$. Given $\bm{x},\bm{y}\in\mathbb{R}^{d}$, the notation $[\bm{x}\;\bm{y}]\in\mathbb{R}^{d\times 2}$ represents the row-wise concatenation, while $[\bm{x};\bm{y}]\in\mathbb{R}^{2d}$ represents the column-wise concatenation.
@@
[…]
@@
^${\hat{\cal{L}}_{\mathcal{S}_{\text{TT}}}}(\bm{W})=\sum_{j=1}^{k}(y_{j}-\text{SM}({\bm{Z}_{j}},{\bm{W}}))^{2}.$ (2)
One can thus refine $\bm{W}^{*}$ by optimizing the above test-time empirical loss. In this paper, we focus on a TTT strategy involving a single gradient descent step over ^${\hat{\cal{L}}_{\mathcal{S}_{\text{TT}}}}$, i.e. ${\bm{W}_{\text{TT}}}:=\bm{W}^{*}-\eta\nabla{\hat{\cal{L}}_{\mathcal{S}_{\text{TT}}}}(\bm{W}^{*})$. The error incurred by the model can then be calculated using the population loss via
@@
where $[\cdot]_{n+1,d+1}$ denotes the entry $(n+1,d+1)$ of the corresponding matrix. Here, the query, key, and value matrices $\bm{W}_{Q},\bm{W}_{K},\bm{W}_{V}\in\mathbb{R}^{(d+1)\times(d+1)}$ are defined as
@@
Here, we collapse the query and key matrices by identifying the top-left $d\times d$ block of $\bm{W}_{Q}\bm{W}_{K}^{\top}\in\mathbb{R}^{(d+1)\times(d+1)}$ with a single matrix $\bm{W}\in\mathbb{R}^{d\times d}$, and choose the value matrix to retrieve the output of the query token with one-layer linear attention. We note that similar models have been used in prior work (Zhang et al., 2023; Mahankali et al., 2023; Ahn et al., 2023a; Li et al., 2024) for the theoretical analysis of a variety of phenomena, albeit not for characterizing the benefits of TTT.
Data model. We consider a linear model with Gaussian data. Specifically, during pre-training, we assume the context tokens ${\bm{z}}_{i}=[\bm{x}_{i};y_{i}]$ and the query input/output vector $[\bm{x};y]$ to be sampled i.i.d. from the distribution ${\cal{D}_{\text{z}}^{\text{PT}}}(\bm{\Sigma}_{\bm{\beta}})$. Concretely, the outputs are generated according to the linear model $y_{i}=\bm{x}_{i}^{\top}{\bm{\beta}_{\text{PT}}}+\xi_{i}$, with task parameter ${\bm{\beta}_{\text{PT}}}\sim\mathcal{N}(\bm{0},\bm{\Sigma}_{\bm{\beta}})$, input features $\bm{x}_{i}\sim\mathcal{N}(\bm{0},\bm{\Sigma}_{\bm{x}})$, and noise terms $\xi_{i}\sim\mathcal{N}(0,\sigma^{2})$ for $i\in[n]$.
@@
[…]
@@
${\bm{X}_{\text{context}}}=\begin{bmatrix}\bm{\bar{x}}_{1}\ \ldots\ \bm{\bar{x}}_{n}\end{bmatrix}^{\top},\quad$ ${\bm{y}_{\text{context}}}=\begin{bmatrix}\bar{y}_{1}~{}\dots~{}\bar{y}_{n}\end{bmatrix}^{\top},$
${\bm{X}_{\text{train}}}=\begin{bmatrix}\bm{\bar{x}}_{n+1}\ \ldots\ \bm{\bar{x}}_{n+k}\end{bmatrix}^{\top},\quad$ ${\bm{y}_{\text{train}}}=\begin{bmatrix}\bar{y}_{n+1}~{}\dots~{}\bar{y}_{n+k}\end{bmatrix}^{\top}.$
@@ 4 Analysis for Isotropic Covariances
Figure 1: Plot of the normalized population losses after test-time training when $\bm{\Sigma}_{\bm{x}}=\bm{\Sigma}_{\bm{\beta}}={\bm{I}}$, $\sigma^{2}=0$, and the new task is ${\bm{\beta}_{\text{TT}}}=\bm{1}_{d}$. Solid lines denote theoretical predictions which match the empirical (markers) results. (a): The figure illustrates a non-monotonic trend as the ratio $\tfrac{n}{d}$ shifts, see Corollary 4.3. Setting: $n=300$; the ratio $\frac{n}{d}$ changes between $0.1$ and $4$; the three lines depict $\gamma=1$, $\gamma=2$, and $\gamma=100$ where $\gamma=k/d$; the step size $\eta$ is selected optimally as per Theorem 4.2. (b): The figure reveals a threshold in $k$ at which the preferable initialization switches from the optimal pre-trained $\bm{W}^{*}$ to the null model $\bm{0}_{d\times d}$ as $k$ increases. The transition threshold $k$ aligns well with Corollary 4.5. Setting: $n=200$; $d=400$; $k$ changing between 00 and $4n$ with equal increments.
@@
Given $k,n,$ and $d$, the loss of the pre-trained parameters, ^${\hat{\cal{L}}_{\mathcal{S}_{\text{TT}}}}(\bm{W}^{*})$, is a random variable with respect to the test-time set $\mathcal{S}_{\text{TT}}$. Consequently, the loss ${\cal{L}}({\bm{W}_{\text{TT}}})$ is also a random variable dependent on $\mathcal{S}_{\text{TT}}$. When applying a single-step update to the pre-trained model’s parameters, we define the expectation of the loss ${\cal{L}}({\bm{W}_{\text{TT}}})$ with respect to $\mathcal{S}_{\text{TT}}$ in (4), treating it as a function of the step size $\eta$. The optimal step size is then selected to minimize the expected test-time training loss. In the following theorem (proved in Appendix B), we present the optimal learning rate and the corresponding improvement in the loss induced by test-time training.
@@
The approximations above are robust, and the expressions become exact as $n,d\to\infty$ while keeping the ratio $n/d$ fixed; see Remark B.4. Theorem 4.2 establishes the improvement produced by the test-time training as a function of $k,n,$ and $d$. When $n$ and $d$ are fixed, such improvement is proportional to $\frac{k}{k+d}$. Additionally, if $k/d=\gamma$ is fixed, the loss ${\cal{L}}_{TT}({\bm{W}_{\text{TT}}})$ exhibits an intriguing behavior as a function of $n/d=\alpha$, as described by the next corollary (proved in Appendix B).
@@
The phase transition points described in Corollary 4.3 are approximate but become exact as $k,n,$ and $d$ approach infinity while maintaining the ratios $\gamma=k/d$ and $\alpha=n/d$.
@@
We provide the proof of Theorem 4.4 in Appendix B, where we also solve the more general noisy setting with initial weight $\bm{W}^{*}=\bm{0}_{d\times d}$. If $\sigma^{2}=\mathcal{O}\left(\|{\bm{\beta}_{\text{TT}}}\|_{2}^{2}\right)$ and $k$ grows sufficiently faster than $d$ (e.g., $k/d\to\infty$), then the test-time training update can reduce the loss to near 00 as $k,n,d\to\infty$ with $\alpha=n/d=\Theta(1)$.
Armed with Theorem 4.2 and 4.4, we can now establish when it is better to initialize with the pre-trained $\bm{W}^{*}$, as opposed to the zero (null) initialization, as the size $k$ of the test-time training set varies.

### v2

@@ 1 Introduction
Modern language models can be viewed as general-purpose, addressing a diverse range of user queries in a zero-shot fashion Kojima et al. (2022); Wei et al. (2022). However, as the complexity or novelty of the query increases, such as in complex multi-step reasoning scenarios, the pretrained model may falter. This has motivated two popular approaches: in-context learning and test-time computation. In-context learning (ICL) incorporates demonstrations related to the query as part of the prompt facilitating better inference by the model Brown et al. (2020); Min et al. (2022). Test-time computation methods explicitly increase the inference-time compute to elicit higher quality responses Snell et al. (2024); Jaech et al. (2024). An important instance of test-time computation is test-time training (TTT) where the weights of a model are explicitly updated to adapt to specific test instances Sun et al. (2020); Liu et al. (2021). A gradient-based TTT approach can be described as follows: Given a test prompt and a pretrained sequence model, we update the pretrained weights by performing a few gradient iterations on a suitable self-supervised objective (e.g. next-token prediction objective). We can then use the resulting model to perform the inference on the test prompt. This procedure is applicable beyond language models, and it is conceptually similar to meta-learning approaches such as model-agnostic meta-learning Finn et al. (2017).
Recently, TTT has found significant success in the context of language modeling by boosting the accuracy and reasoning capability of state-of-the-art models Akyürek et al. (2024); Sun et al. (2024). This success is in part due to the fact that TTT can be naturally integrated within in-context learning: We can fine-tune the transformer model to fit to the labels or chain-of-thought rationales provided as part of the in-context demonstrations, and then re-apply ICL on the adapted model. In fact, recent work Akyürek et al. (2024) utilizes a variation of this procedure and additional data augmentation to obtain remarkable improvements in the ARC reasoning benchmark. This motivates the central question of our work:
@@
Contributions. As our central contribution, we address this question by providing a comprehensive theoretical study of test-time training for one-layer linear transformers. Focusing on prompts following a linear dataset model, we provide a precise risk characterization of TTT with one gradient step update rule. This characterization is established in terms of three ingredients: (i) context length (number of in-context examples during inference), (ii) target sample size available for TTT in Section 4, and (iii) alignment between the pre-trained model and target task in Section 5. We show that, as sample size increases, TTT can alleviate the distribution shift bottlenecks that arise in standard ICL. This also reveals regimes where TTT with zero or small initialization is preferable to TTT with the pre-trained model (i.e. cold vs. warm start). Interestingly, our experiments with the GPT2 architecture Radford et al. (2019) show that multilayer transformers exhibit behavior in line with our theory. Our theory and experiments demonstrate that one step of TTT yields significant performance gains with less computation, consistent with recent empirical observations (Akyürek et al., 2024) that only a few gradient steps offer substantial test-time improvements. Additionally, while standard ICL requires $\Omega(d)$ context length under an isotropic task prior, with $d$ denoting feature dimension, we prove that TTT can succeed with $o(d)$ context length by effectively memorizing the target task. Our technical novelty stems from accurately capturing the statistical benefit of TTT during in-context learning. Specifically, while the transformer is trained on a single prompt, we characterize how the sample complexity benefit of TTT is proportional to the number of target examples within the prompt.
As empirical corroboration of our theory, we explore tabular learning and TabPFN Hollmann et al. (2023, 2025) – a state-of-the-art tabular foundation model pretrained with structural causal model priors. TabPFN is well-aligned with our theoretical setting with similar token encodings but different prior distributions. An important drawback of TabPFN lies in its inference cost, as it uses a full tabular dataset as context during inference. In line with our theory, we demonstrate that TTT can convert TabPFN into a task-specific tabular model that works equally well with significantly less data (up to 5 times fewer). This in turn implies substantial inference gains given that the complexity of softmax-attention is quadratic in the sequence length.
@@ 2 Related Work
In-context learning. In-context learning has received significant interest during the past few years (Brown et al., 2020; Liu et al., 2023; Agarwal et al., 2024). This has also motivated works toward a stronger theoretical understanding of ICL (Zhang et al., 2024; Mahankali et al., 2024; Ahn et al., 2023b; Xie et al., 2022; Garg et al., 2022; Li et al., 2023). Closer to us, Mahankali et al. (2024); Ahn et al. (2023b); Zhang et al. (2024) study the optimization landscape of a one-layer linear attention model and show that the optimized model implements a single step of projected gradient descent over the in-context demonstrations. More recent works Lu et al. (2024); Wu et al. (2023) extend these to characterize the pretraining task sample complexity of ICL. While these works focus on pretraining capabilities, we focus on the adaptation of the pretrained model to the target task. Additionally, rather than empirical risk minimization, we use gradient descent for adaptation and characterize its risk when the sample size is determined by the target prompt length. In our theoretical model, each token represents a data point containing input features and the target label. Remarkably, TabPFN Hollmann et al. (2023, 2025) shows that, using this simple encoding and pretraining the model with sufficiently rich data priors, transformers can accomplish state-of-the-art classification on tabular datasets via in-context learning. There have been also efforts (Thomas et al., 2024) to fine-tune in-context tabular models like TabPFN at the dataset level using retrieval-based local context selection.
Test-time training. Test-time training Sun et al. (2020); Liu et al. (2021); Gandelsman et al. (2022) and related test-time adaptation Wang et al. (2021); Niu et al. (2022); Yuan et al. (2023) methods aim to overcome distribution shift bottlenecks. This is typically accomplished by adapting the model with the test example using self-supervised or unsupervised objectives. These methods can work with just a single text example or admit a minibatch of examples. For sequence/language modeling tasks, one can utilize TTT on the test sequence via the next-token prediction objective Sun et al. (2024); Hardt and Sun (2024); hübotter2025efficientlylearningtesttimeactive. Specifically, during in-context learning, the query is unlabeled (e.g. math problem to solve), but we are provided with related examples and associated labels (e.g. through retrieval). Thus, we can fit to these labels as the source of supervision (essentially fine-tuning the model on the dataset of in-context examples). For instance, the training objective in Akyürek et al. (2024) utilizes this approach boosted by additional data augmentation. Finally, the computational efficiency of TTT is an important consideration which motivates our investigation of TTT with a single gradient update.
@@ 3 Problem Setup
Notation. Let $[n]$ denote the set $\{1,\cdots,n\}$ for an integer $n\geq 1$. We denote vectors and matrices using bold lower-case and upper-case letters, respectively, such as $\bm{x}$ for vectors and $\bm{X}$} for matrices. The element $x_{i}$ refers to the $i$-th entry of the vector $\bm{x}$. We represent the zero vector of size $n$ by $0_{n}$ and the zero matrix of size $m\times n$ by $\bm{0}_{m\times n}$. The operator $\texttt{tr}({\bm{X}})$ represents the trace of $\bm{X}$}, $\bm{X}$$^{\dagger}$ is its Moore–Penrose pseudoinverse, and $\|{\bm{X}}\|_{2}$ denotes the spectral norm of $\bm{X}$}. Given $\bm{x},\bm{y}\in\mathbb{R}^{d}$, the notation $\bm{x}\;\bm{y}]\in\mathbb{R}^{d\times 2}$ represents the row-wise concatenation, while $\bm{x};\bm{y}]\in\mathbb{R}^{2d}$ represents the column-wise concatenation.
@@
To estimate the output query $y$, we focus on a sequence model $\text{SM}(\bm{Z},\bm{W})$ with $\bm{Z}$ the input prompt and $\bm{W}$ the model parameters. We assume that the sequence model is pre-trained with a token distribution $\cal{D}_{\text{z}}^{\text{PT}}$} where $({\bm{z}}_{i})_{i=1}^{n},[\bm{x};y]\overset{\text{i.i.d.}}{\sim}{\cal{D}_{\text{z}}^{\text{PT}}}$ with the corresponding optimal parameters given by
@@
${\hat{\cal{L}}_{\mathcal{S}_{\text{TT}}}}(\bm{W})=\sum_{j=1}^{k}(y_{j}-\text{SM}({\bm{Z}_{j}},{\bm{W}}))^{2}.$ (2)
One can thus refine $\bm{W}^{*}$ by optimizing the above test-time empirical loss. In this paper, we focus on a TTT strategy involving a single gradient descent step over $\hat{\cal{L}}_{\mathcal{S}_{\text{TT}}}$}, i.e. ${\bm{W}_{\text{TT}}}:=\bm{W}^{*}-\eta\nabla{\hat{\cal{L}}_{\mathcal{S}_{\text{TT}}}}(\bm{W}^{*})$. The error incurred by the model can then be calculated using the population loss via
@@
where $\cdot]_{n+1,d+1}$ denotes the entry $(n+1,d+1)$ of the corresponding matrix. Here, the query, key, and value matrices $\bm{W}_{Q},\bm{W}_{K},\bm{W}_{V}\in\mathbb{R}^{(d+1)\times(d+1)}$ are defined as
@@
Here, we collapse the query and key matrices by identifying the top-left $d\times d$ block of $\bm{W}_{Q}\bm{W}_{K}^{\top}\in\mathbb{R}^{(d+1)\times(d+1)}$ with a single matrix $\bm{W}\in\mathbb{R}^{d\times d}$, and choose the value matrix to retrieve the output of the query token with one-layer linear attention. We note that similar models have been used in prior work Zhang et al. (2023); Mahankali et al. (2023); Ahn et al. (2023a); Li et al. (2024, 2025) for the theoretical analysis of a variety of phenomena, albeit not for characterizing the benefits of TTT.
Data model. We consider a linear model with Gaussian data. Specifically, during pre-training, we assume the context tokens ${\bm{z}}_{i}=[\bm{x}_{i};y_{i}]$ and the query input/output vector $\bm{x};y]$ to be sampled i.i.d. from the distribution ${\cal{D}_{\text{z}}^{\text{PT}}}(\bm{\Sigma}_{\bm{\beta}})$. Concretely, the outputs are generated according to the linear model $y_{i}=\bm{x}_{i}^{\top}{\bm{\beta}_{\text{PT}}}+\xi_{i}$, with task parameter ${\bm{\beta}_{\text{PT}}}\sim\mathcal{N}(\bm{0},\bm{\Sigma}_{\bm{\beta}})$, input features $\bm{x}_{i}\sim\mathcal{N}(\bm{0},\bm{\Sigma}_{\bm{x}})$, and noise terms $\xi_{i}\sim\mathcal{N}(0,\sigma^{2})$ for $i\in[n]$.
@@
We construct the test-time set $\mathcal{S}_{\text{TT}}=\{(\bm{Z}_{j},y_{j})\}_{j=1}^{k}$ by following the test-time distribution ${\cal{D}_{\text{z}}^{\text{TT}}}({\bm{\beta}_{\text{TT}}})$. We first sample $(n+k)$ query input/output pairs $\{(\bm{\bar{x}}_{i},\bar{y}_{i})\}_{i=1}^{(n+k)}\overset{\text{i.i.d.}}{\sim}{\cal{D}_{\text{z}}^{\text{TT}}}({\bm{\beta}_{\text{TT}}})$. Then, we designate the first $n$ samples as fixed context tokens across the set, whereas we use the latter $k$ samples as queries in the set. In formulas,
@@
${\bm{X}_{\text{context}}}=\begin{bmatrix}\bm{\bar{x}}_{1}\ \ldots\ \bm{\bar{x}}_{n}\end{bmatrix}^{\top}$$\quad$ ${\bm{y}_{\text{context}}}=\begin{bmatrix}\bar{y}_{1}~\dots~\bar{y}_{n}\end{bmatrix}^{\top}$,
${\bm{X}_{\text{train}}}=\begin{bmatrix}\bm{\bar{x}}_{n+1}\ \ldots\ \bm{\bar{x}}_{n+k}\end{bmatrix}^{\top}$$\quad$ ${\bm{y}_{\text{train}}}=\begin{bmatrix}\bar{y}_{n+1}~\dots~\bar{y}_{n+k}\end{bmatrix}^{\top}$.
@@ 4 Analysis for Isotropic Covariances
Figure 1: Plot of the normalized population losses after test-time training when $\bm{\Sigma}_{\bm{x}}=\bm{\Sigma}_{\bm{\beta}}={\bm{I}}$, $\sigma^{2}=0$, and the new task is ${\bm{\beta}_{\text{TT}}}=\bm{1}_{d}$. Solid lines denote theoretical predictions which match the empirical (markers) results. (a): The figure illustrates a non-monotonic trend as the ratio $\tfrac{n}{d}$ shifts, see Corollary 4.4. Setting: $n=300$; the ratio $\frac{n}{d}$ changes between $0.1$ and 44; the three lines depict $\gamma=1$, $\gamma=2$, and $\gamma=100$ where $\gamma=k/d$; the step size $\eta$ is selected optimally as per Theorem 4.2. (b): The figure reveals a threshold in $k$ at which the preferable initialization switches from the optimal pre-trained $\bm{W}^{*}$ to the null model $\bm{0}_{d\times d}$ as $k$ increases. The transition threshold $k$ aligns well with Corollary 4.6. Setting: $n=200$; $d=400$; $k$ changing between 0 and $4n$ with equal increments.
@@
Given $k,n,$ and $d$, the loss of the pre-trained parameters, ${\hat{\cal{L}}_{\mathcal{S}_{\text{TT}}}}(\bm{W}^{*})$, is a random variable with respect to the test-time set $\mathcal{S}_{\text{TT}}$. Consequently, the loss ${\cal{L}}({\bm{W}_{\text{TT}}})$ is also a random variable dependent on $\mathcal{S}_{\text{TT}}$. When applying a single-step update to the pre-trained model’s parameters, we define the expectation of the loss ${\cal{L}}({\bm{W}_{\text{TT}}})$ with respect to $\mathcal{S}_{\text{TT}}$ in (4), treating it as a function of the step size $\eta$. The optimal step size is then selected to minimize the expected test-time training loss. In the following theorem (proved in Appendix B), we present the optimal learning rate and the corresponding improvement in the loss induced by test-time training.
@@
Remark 4.3.
We note that in the proportional $n,d$ regime, the Gaussian approximation in Lemma B.1, which is utilized in the proof of Theorem˜4.2, introduces an error of $\mathcal{O}(n^{-1})$. Additionally, during the derivation, we omit lower-order terms (such as $2n$ in $n^{2}+2n$), which introduces errors of one degree lower relative to the leading terms. Consequently, after carrying these approximations through the entire derivation, the overall approximation error remains at most $\mathcal{O}(n^{-1})$, since the final loss expression is zeroth-order in $n,d$. Hence, our result is robust to these errors as $n,d\rightarrow\infty$ while maintaining the ratio $\alpha=n/d$. The same reasoning also applies to the non-isotropic covariance case, which will be discussed in Section˜5.
Theorem 4.2 establishes the improvement produced by the test-time training as a function of $k,n,$ and $d$. When $n$ and $d$ are fixed, such improvement is proportional to $\frac{k}{k+d}$. Additionally, if $k/d=\gamma$ is fixed, the loss ${\cal{L}}_{TT}({\bm{W}_{\text{TT}}})$ exhibits an intriguing behavior as a function of $n/d=\alpha$, as described by the next corollary (proved in Appendix B).
@@
The phase transition points described in Corollary 4.4 are approximate but become exact as $k,n,$ and $d$ approach infinity while maintaining the ratios $\gamma=k/d$ and $\alpha=n/d$. The non-monotonic behavior identified in Corollary˜4.4 is observed as a result of two opposing effects, which are the initial loss and the improvement by TTT. As $n$ grows against $d$ (i.e. $\alpha$ increases), the pre-trained model does better initially and already has a lower loss before TTT, which makes it harder to be further reduced by TTT as the rank-1 update is unable to correct all directions. On the other hand, when $d$ grows against $n$ ($\alpha$ decreases), the initial loss is high, providing more room (error) to be corrected by rank-1 update, and thus, the improvement by TTT is larger. This intuition aligns with Theorem 4.2, which establishes that the TTT improvement scales as $(\frac{d}{n+d})^{3}$. Together, these two trends result in the non-monotonic behavior.
@@
We provide the proof of Theorem 4.5 in Appendix B, where we also solve the more general noisy setting with initial weight $\bm{W}^{*}=\bm{0}_{d\times d}$. If $\sigma^{2}=\mathcal{O}\left(\|{\bm{\beta}_{\text{TT}}}\|_{2}^{2}\right)$ and $k$ grows sufficiently faster than $d$ (e.g., $k/d\to\infty$), then the test-time training update can reduce the loss to near 0 as $k,n,d\to\infty$ with $\alpha=n/d=\Theta(1)$.
Armed with Theorem 4.2 and 4.5, we can now establish when it is better to initialize with the pre-trained $\bm{W}^{*}$, as opposed to the zero (null) initialization, as the size $k$ of the test-time training set varies.